How do you make the following 300 mL solution:

10 mM Tris

25 mM NaCl

20% Sucrose

3 ug/mL BSA

.72 M KCL

Use these as your givens:

2 M Tris

Solid NaCl (MW = 58.44 g/mol)

Solid Sucrose (MW = 342.3 g/mol)

3 mg/mL BSA

Solid KCL (MW = 74.5513 g/mol)

Thanks!
Answer #1

Answer:

(i) Preparation of 300 mL 10 mM Tris solution:

Supplied, 2 M Tris solution

We know that, V1S1 = V2S2

where, V1 = volume of Tris when strength S1 = 2M = 2000 mM, V2 = 300 mL, S2 = 10 mM

Therefore, V1 × 2000 mM = 300 mL × 10 mM

∴ V1 = (300 × 10) / 2000 mL = 1.5 mL

Thus, 1.5 mL of 2M Tris must be added to 298.5 mL water to make 300 mL 10 nM Tris solution.

(ii) Preparation of 300 mL 25 mM NaCl solution:

Given, MW of NaCl = 58.44 g/mol

We know that, for 1000 mL of 1 mM NaCl solution amount of NaCl needed = 58.44 mg

Therefore, for 300 mL of 25 mM NaCl solution amount of NaCl needed

= (58.44 × 300 × 25) / 1000 mg = 438.3 mg

Thus, 438.3 mg NaCl must be added to 300 mL water to make 300 mL of 25 mM NaCl solution.

(iii) Preparation of 300 mL 20% Sucrose solution:

Given, MW of Sucrose = 342.3 g/mol

We know that, for 100 mL of 1% sucrose solution amount of sucrose needed = 1 g

Therefore, for 100 mL of 20% sucrose solution amount of sucrose needed = 20 g

Therefore, for 300 mL of 20% sucrose solution amount of sucrose needed

= (300 × 20) / 100 g = 60g in 300 mL water.

(iv) Preparation of 300 mL 3 ug/mL BSA solution:

Supplied, 3 mg/mL BSA ≡ 3000 ug/mL

Therefore, for 300 mL of 3 ug/mL BSA solution amount of supplied BSA solution needed

= (300 × 3) / 3000 ug/mL = 0.3 mL

Therefore, for 300 mL of 3 ug/mL BSA solution amount of supplied BSA solution needed 0.3 mL and rest water.

(v) Preparation of 300 mL 0.72 M KCl solution:

Given, MW of KCl = 74.5513 g/mol

We know that, for 1000 mL of 1 M KCl solution amount of KCl needed = 74.5513 g

Therefore, for 300 mL of 0.72 M KCl solution amount of KCl needed

= (74.5513 × 300 × 0.72) / 1000 g = 16.103 g

Thus, 16.103 g KCl must be added to 300 mL water to make 300 mL of 0.72 M KCl solution.